OpenStudy (anonymous):

all degree solutions for: cos 4x=-1

OpenStudy (anonymous):

try u substitution. set u=4x, and solve like a normal equation. then substitute 4x back in

OpenStudy (anonymous):

cos4x= 8cos^4(x) - 8cos^2(x) + 1 now 8cos^4(x) - 8cos^2(x) + 1=-1 cos^2x= a 8a^2-8a +1 =-1 8a^2 -8a +2 =0 a=1/2 since its a Sqr root answers will be +sqr 1/2 and -sqr 1/2 angle will be = 60 deg now from unit circle Cos is positive in 1st and last quadrants hence answers will be = 60 deg and 300 deg cos is negative in 2nd and 3rd quadrants hence answers will be = 120 deg and 240 deg

OpenStudy (sirm3d):

the general formula for the solution when the coefficient a is other than 1 or -1 is \[ax=x_i+n(360)\] for integers n=0,1,2,..., a-1

OpenStudy (anonymous):

is it cos^4(x) or cos(4x)??

OpenStudy (sirm3d):

where \[x_i\] is a solution you get via a calculator (or inspection) for the problem, \[x_i=180\]

OpenStudy (sirm3d):

ugh. \[\cos x=\frac{ \sqrt{2} }{ 2 }\] so \[x=45^o\] and not 60 degrees.

OpenStudy (anonymous):

well done sirm! so lets correct the answers 45 and 315 135 and 225